# Towards Autonomous Driving with Small-Scale Cars: A Survey of Recent Development

### v1

@@ III-G3 Drifting
III-H Cooperative driving
In comparison to the scenarios mentioned before which typically only involve one controlled vehicle, cooperative driving refers to a collaborative approach where multiple autonomous vehicles interact and communicate with each other to achieve common goals, navigate shared spaces, and enhance overall traffic efficiency and safety. It relies on vehicles exchanging information, such as their positions, speed, intended trajectories, and other relevant data, to make collective decisions. The decision-making processes of cooperative driving include centralized control and decentralized control. In the centralized control system, a central controller oversees and coordinates the actions of all vehicles involved in the cooperative driving system. This central entity collects data from all vehicles, processes the information, and makes decisions. On the contrary, in a decentralized control system, decision-making is distributed among individual vehicles within the environment. Each autonomous vehicle makes independent decisions based on local information gathered from its sensors and communication with nearby vehicles. The vehicles negotiate and collaborate among themselves, sharing relevant data and coordinating actions to achieve common goals.
Figure 12: Overall schematics of the multi-vehicle, mixed reality RL approach in [41]. Both virtual and real DeepRacer vehicles exist within the simulation that manages the physics of the virtual cars and emulates collisions in mixed reality.
Examples of cooperative driving encompass various aspects, spanning from mapping, path planning, obstacle avoidance to roundabout navigation, and traffic jam mitigation. For example, in the study by [237], a feature-based map registration method is utilized for a collaborative SLAM system. Three distinct vehicles collaborate, enhancing SLAM efficiency by achieving faster mapping capabilities through collaborative efforts. In the work by [238], a multi-vehicle path planning strategy Enhanced Conflict Based Search with Task Assignment (ECBS-TA) [239] is employed for MuSHR cars. It involves optimal task assignment considering discrete workspace representation. Following this, collision-free trajectories are planned for all vehicles using decentralized MPC for precise path tracking. Addressing safety concerns, [240] devises a distributed priority assignment algorithm called priority-based non-cooperative distributed model predictive control (P-DMPC). This algorithm elevates the priority of a vehicle based on the number of potential collisions anticipated along its planned trajectory, thereby enhancing safety in multi-vehicle scenarios. [41] introduces a mixed reality framework depicted in Fig. 12 tailored for safe and efficient RL driving policies in multi-vehicle systems. This setup includes one physical vehicle and sixteen virtual vehicles. The RL agent demonstrates capabilities in lane keeping and obstacle avoidance within this complex multi-vehicle environment. In the case of navigation through a roundabout, the vehicles employ cooperative strategies to merge into the appropriate lane within the roundabout by communicating their intentions to enter the roundabout and negotiate their entry sequence. Vehicles need to follow certain merge and yield protocols, which for example vehicles yield to others already within the roundabout, vehicles inside the roundabout adjust their speeds and trajectories. To achieve this scenario, [59] utilizes RL algorithms to control a set of miniature cars successfully coordinated at a roundabout and ensure a smooth merge, and demonstrates the results in University of Delaware’s Scaled Smart City (UDSSC) testbed. Later, [57] develops a decentralized optimal control framework for 9 cars in a multi-lane roundabout in UDSSC and eliminates stop-and-go driving while preserving safety.
Another driving scenario to discuss is traffic jam mitigation, which remains a focal point in cooperative driving, emphasizing the collaborative endeavors among multiple autonomous vehicles to prevent or alleviate traffic congestion. The primary goal is to sustain optimal traffic flow and speed, particularly in unexpected scenarios such as abrupt braking or lane closures. This requires vehicles to coordinate actions collectively and respond to dynamic environmental changes. For instance, [66] introduces a framework that modifies the intelligent driver model (IDM) [241] and the MOBIL lane change model [242] for cooperative driving. Additionally, [243] and [244] propose decentralized optimal traffic control methods. Their approaches demonstrate reduced travel times compared to baseline scenarios involving human-driven vehicles lacking connectivity. [24] presents the CoRoLa Car Platform, and achieves multi-vehicle platooning driving with Adaptive cruise control (ACC). Furthermore, recent research has explored human-robot cooperation ([245, 246]), as well as collaborations between Unmanned Ground Vehicles (UGVs) and Unmanned Aerial Vehicles (UAVs) [247]. These studies delve into cooperative methodologies between different vehicle types, fostering improved traffic management and congestion mitigation through collaborative efforts.
Figure 13: Diagram of trajectory planner with C-MOBIL and C-IDM in [66]. The ego vehicle plans a trajectory and velocity profile based on its state and the actual state of the neighboring vehicles.
TABLE II: Overview of research in different tasks for small-scale cars.
Task Category Method References
Localization and Mapping KF [108, 109, 110, 219]: BARC; [111]: QCar
PF [45]: F1TENTH; [114]: RACECAR
GMapping [122]: Roborace; [123, 125]: -
HectorSLAM [119, 120, 121]: -
Cartographer [248]: F1TENTH; [127]: Innopolis UGV; [126, 128]: -
GALNet [133]: Autominy
PACT [134]: MuSHR
LoFTR [135]: Jetbot
Path planning Classical Dijkstra’s algorithm [142]: -
Hybrid A star [143]: QCar
GA [145, 146]: -
RRT [147]: RACECAR
ML-based RL [149]: Jetbot; [150]: MuSHR; [98]: Duckiebot; [148]: -
BRPO [151]: MuSHR
IL [249]: Donkeycar
Path following Kinematic controller [154]: QCar; [152, 153]: -
MPC [158]: MuSHR; [155, 156, 157]: -
IL [159]: ART/ATK
DRL [160]: Autominy; [161]: JetRacer
Lane keeping Classical classical CV + PID [7, 250]: Duckiebot
classical CV + rule-based [164]: Donkeycar
ML-based CNN + PID [166]: Duckiebot
classical CV + RL [95, 94]: Duckiebot
CNN + RL [43]: Donkeycar
ViT [168]: Duckiebot
IL (raw input) [49, 173]: DeepRacer; [92, 174]: Donkeycar; [176]: Duckiebot
RL (raw input) [10, 177, 178, 181, 38]: DeepRacer; [179, 99]: Duckiebot; [180]: JetRacer; [47]: F1TENTH
IL (autoencoder) [183, 184]: Duckiebot; [251]: -
RL (autoencoder) [42]: Donkeycar; [185]: Duckiebot
Obstacle Detection Vision-based Appearance-based [36]: Duckiebot
Motion-based [195, 196]: -
Depth-based [198, 197]: -
CNN [65]: Go-CHART; [203]: JetRacer; [204]: Duckiebot; [252]: Jetbot; [253, 254]: -
LiDAR-based Minimum value [48, 206]: F1TENTH
FTG [12]: F1TENTH
Racing Classical iSAM2 + MPC [9, 51]: Autorally
Particle filter + MPC [55, 53]: Autorally; [218]: F1TENTH
CNN + TPVDCM [214]: Autorally
MPC [15, 255, 256]: ORAC
LMPC [217, 219]: BARC
LPV-MPC [110]: BARC
PAPC [54]: Autorally
Pure pursuit [220]: TUNERCAR
End-to-end IL [91]: Autorally
BNN [225]: Autorally
RL [226, 227]: F1TENTH; [222]: ORAC; [228, 229, 232, 224]: -
Drifting EKF, LQR [8, 234]: BARC
NMPC [235]: BARC; [236]: RACECAR
Car following DRL [95]: Duckiebot; [187]: Donkeycar; [188]: QCar
Overtaking DRL [94]: Duckiebot
MPC [190]: MuSHR
Traffic sign recognition Classical [37]: -
Tiny-YOLO [210]: -
Cooperative driving Centralized control [41]: DeepRacer; [59]: UDSSC; [24]: CoRoLa; [246]: Jetbot; [247]: QCar
Decentralized control [237]: QCar; [238]: MuSHR; [240]: $\mu$Car; [57, 243, 244, 16]: UDSSC; [66]: CamMini; [245]: Jetbot
IV Sensors and Sensor systems
In Section II-A, various platforms are explored, each employing diverse sensors to perceive their surroundings and execute the tasks detailed in Section III. Following that, we discuss the most commonly utilized sensors across these platforms, providing their respective use cases.
IV-A Camera
An RGB camera provides a stream of information that is easily processed and understood by humans. This helps researchers to better understand the perspective of the robot. The amount of information provided by cameras is huge, which allows the robot to perceive a lot of information with just one sensor in a short amount of time. The downside of this is, that they need a high bandwidth communication link to whichever part of the robot needs those information. Also with the information in images being so densely encoded, it is difficult for a robot to understand the environment.
The field of computer vision has emerged as a crucial scientific discipline, facilitating the extraction of pertinent details from images and videos. In recent years, the advent of deep learning models, such as YOLO [199], has significantly contributed to enhancing the robot’s ability to interpret scenes captured by cameras. It is worth noting that cameras face external influences, such as varying lighting conditions, which can substantially alter the captured images. In extreme scenarios, such as very dark environments or instances where direct light affects the camera, the effectiveness of cameras may be compromised. A big advantage of cameras is their wide pricing range, starting from 15 USD, corresponding to images of different quality in regards to resolution and dynamic range. They are also easy to interface with, usually requiring only a USB connection. Some tasks can be accomplished only with cameras, e.g. traffic sign or road marking detection [210, 257, 258].
Beyond RGB cameras, more advanced options exist, including those detecting optical flow [259] or depth cameras [260], employing different methods to ascertain the distance to objects. These advanced cameras provide both depth information and standard RGB images. In the context of autonomous tasks for single-car platforms discussed in this paper, cameras emerge as a ubiquitous choice. Their cost-effectiveness and information-rich output make them the preferred sensor for many platforms. Among the commonly used cameras in scaled vehicles, the Raspberry Pi Camera stands out [7, 11, 31, 61, 22, 73, 29, 30, 65, 24, 60]. Designed for the Raspberry Pi, it combines affordability with decent image quality, offering a resolution of 1080p and a dynamic range of 44dB. Another popular option used in the discussed platforms are the Intel Realsense [^4] https://www.intelrealsense.com/ line of cameras [19, 261, 21]. Those include two cameras and an IR projector, which allow to capture not only an RGB image but also depth information of each pixel by means of stereovision. The IR projector helps to improve accuracy in scenes with poor textures. The manufacturer claims an accuracy in the centimeter range.
IV-B LiDAR
LiDAR, another widely employed sensor in numerous platforms, functions by measuring the time taken for a beam of light to reflect off a surface. This beam is emitted from multiple angles (typically ranging from 360 to 2048) around the LiDAR, creating a detailed map of points representing objects in the platform’s vicinity. There are also multi-layer LiDARs with laser beams spanning up to 128 angles, providing the platform with information about objects at different heights and generating a comprehensive 3D point cloud. Available in various resolutions, LiDARs offer enhanced information, particularly in intricate and structured environments. The typical scanning frequency of LiDARs ranges from 5 to 20 Hz, with high-end models reaching a detection range of up to 350 meters.
The distinct advantage of LiDAR over cameras lies in its direct identification of the platform’s surroundings, including the precise position of objects relative to the robot. In contrast, a camera first needs to identify an object in an image and then estimate its position. Moreover, LiDAR provides significantly more accurate positional information compared to estimations derived from camera images. However, unlike cameras, LiDARs lack the ability to discern details such as the color of the detected objects. This limitation renders them unsuitable, for example, in tasks like traffic signs or lane detection. LiDARs also possess a notable advantage over cameras as they remain unaffected by environmental conditions such as varying lighting conditions, and they can operate effectively even in complete darkness. Additionally, certain LiDAR models provide information on reflectivity levels, enabling a rough distinction between different surface types. For instance, they can differentiate highly reflective road markings from less reflective obstacles. LiDARs are typically more expensive than cameras. The cheapest hobby LiDARs start at around 150 USD for a simple one plane sensor. But the prices quickly go up as the resolution or amount of planes increases.
Typically physically larger platforms use LiDAR sensors [10, 11, 12, 18, 8, 19, 20, 21, 261, 262, 65, 23, 60, 13] as they tend to be bigger and heavier than cameras. Those platforms are also usually more versatile and used for more different tasks and therefore the additional cost of a LiDAR is feasible. In our investigation, all platforms utilizing LiDAR also integrate a camera to address the limitations of LiDAR for specific tasks. The LIDAR that is used by far the most on the platforms we investigated is the single plane RPLIDAR  [^5] https://www.slamtec.ai/product/slamtec-rplidar-a3/ [11, 8, 20, 21, 261, 65, 23, 60]. It offers a reasonable detection range of 12m and a resolution of 0.225 degrees for small scale car platforms with a 360 field of view. And is one of the cheaper options costing around 400 USD. But it only allows for a 10 Hz scan frequency. The second most popular, higher end solution is the Hokuyo UST-10LX [12, 18], offering a 270 degree field of view and a 10m detection range.
IV-C IMU and Encoders
The information of these two types of sensors is usually combined with other absolute positioning inputs like a camera or a central localization system to fuse them in a filter like a Kalman filter [106] or in any other kind of perception module. IMUs are available at relatively affordable prices, ranging from 15 USD. The two IMU sensors most commonly used in the discussed platforms are the MPU9250 and MPU6050 [7, 11, 5, 21, 261, 60]. Although it has to be noted, that not all publications mention the exact model of the IMU used. Both sensors feature a 3-axis gyroscope, which measures the rotational speed, and a 3-axis accelerometer, which measures linear acceleration. The MPU9250 additionally offers a 3-axis magnetometer to incorporate measurements of the Earth’s magnetic field into orientation sensing.
Odometry encoders can be found in several different forms, ranging from magnetic encoders to optical ones. Their implementation is usually dependent on the platform, as they need to be tightly integrated into the mechanics of the robot. The encoders are usually below 10 USD. Almost all of the platforms we investigated used an IMU and some form of encoder as they are cheap and provide good position information, especially in slow indoor environments. The exception was the smallest platform, primarily designed for swarm robotic investigations [17], where extreme cost considerations led to the exclusion of these sensors.
IV-D GPS
GPS is the least frequently used sensor with model scale platforms [9, 8]. Most of the discussed platforms are designed for indoor use, which eliminates the use of GPS as the signal does not penetrate buildings. However, recognizing the pivotal role GPS plays in full-sized vehicles, many model-scale platforms opt to replace it with indoor central localization systems, as detailed in Section IV-E, to enhance realism. GPS receivers are also comparatively cheap at around 30 USD. However, for those seeking higher precision, the cost can escalate to several hundred USD, especially for the more advanced GPS RTK systems [263]. Such precision may become necessary for model-scale platforms, as conventional GPS systems typically only offer an accuracy of approximately 5 meters [264].
IV-E Central Localization system
In situations where the emphasis is on studying the interaction among multiple vehicles rather than individual vehicle behavior, a central localization system streamlines the task, allowing a more focused approach to other aspects such as obstacle avoidance or overtaking. By decoupling the algorithms for (multi) vehicle behavior from the perception of individual vehicles, central systems enable the study of system dynamics in an idealized ”best case” scenario. In this context, the term ”best case” refers to a situation where a vehicle possesses precise knowledge of its own position and that of surrounding obstacles. However, it is essential to acknowledge that this assumption may not align with the complexities of real-world road situations. Conversely, central systems find utility in simulating GPS positioning within actual vehicles.
One of the cheapest options implemented in the considered platforms involves the use of special markers or tags on each vehicle, which are observed by one or multiple overhead cameras [71]. These markers, based on ArUco [265] technology, are strategically placed on vehicles and key positions or obstacles. Knowing the intrinsic and extrinsic parameters of the cameras, existing libraries like OpenCV  [^6] https://opencv.org/ can be employed to detect these markers and estimate their position and orientation. As discussed in Section IV-A, cameras are relatively inexpensive, and the markers merely need to be printed out. While the resulting localization precision is reasonable [266], it does decline with the distance of the markers from the camera. Mitigating this issue involves the use of multiple cameras, necessitating synchronization and precise positioning of all cameras.
Off-the-shelf solutions for tracking also exist from various manufacturers, such as OptiTrack  [^7] https://optitrack.com/ or VICON  [^8] https://www.vicon.com/. These solutions predominantly rely on high-speed IR cameras and reflective markers on the tracked objects. The systems inherently provide position and orientation information for different marker assemblies (referred to as rigid bodies) without requiring additional processing. Notably, the accuracy of these systems surpasses that of solutions utilizing RGB cameras, with manufacturers often specifying sub-millimeter accuracy. However, the upfront cost is relatively high, typically exceeding 10,000 USD, and in some cases, additional licensing fees for the requisite software may apply [66, 16].
IV-F Other sensors
Several platforms discussed in Section II-A employ additional sensors to either emulate sensors found in full-size autonomous cars or fulfill specific tasks.
A prevalent type of sensor utilized in many platforms [7, 5, 32, 33, 21, 65, 6, 67, 69, 24, 267, 29] is the single-point distance sensor, available as either ultrasonic or infrared. These sensors operate on the time-of-flight principle, measuring the time taken for a wave to bounce back into the sensor. Typically positioned at the front of vehicles, they often mimic radar sensors and play a crucial role in tasks such as car following.
Line following sensors, consisting of an array of light sensors aimed at the ground, are employed in certain platforms [21, 16, 21] to assist visual line following, with a camera serving as a secondary input. While these sensors can easily discern the line from the road, a limitation is their coverage area, which is relatively small and located underneath the car. If the vehicle deviates significantly from the line, the sensor may struggle to rediscover it.
One platform [71] utilizes RFID readers to detect specific points of interest on the road, such as the beginning of intersections or as indicators for turns on a crossroad. Another platform [6] employs cliff sensors, which function as downward-facing distance sensors, indicating if the part of the vehicle with the sensor is suspended over a steep drop. This sensor serves to prevent the vehicle from inadvertently falling off a precipice in the environment.
For platforms designed for use with multiple vehicles [17, 31, 6, 30], infrared (IR) transmitters and receivers are employed. These facilitate cost-effective communication between different robots.
IV-G Compute Units
To process the information from the sensors and compute actuator movement according to their task, the platforms require a dedicated computational unit.
The prevailing choice across numerous platforms [7, 11, 21, 61, 30, 261, 16, 66, 62, 63, 71, 24] is the Raspberry Pi single-board computer (SBC) [^9] https://www.raspberrypi.com/. Built on a System on a Chip (SoC), the Raspberry Pi features a compact physical size but offers modest computational performance, with a 4-core 1.5GHz CPU. Despite its relatively low processing power, the Raspberry Pi is well-suited for the majority of tasks performed by these platforms.
A comparable alternative to the Raspberry Pi is the Nvidia Jetson [^10] https://www.nvidia.com/de-de/autonomous-machines/embedded-systems/. Also based on an SoC, the Nvidia Jetson stands out by incorporating an additional GPU to accelerate machine learning tasks. Given their similar connectivity, some platforms offer compatibility with either the Raspberry Pi or the Nvidia Jetson [7, 11, 261]. However, the majority of platforms exclusively support the Nvidia Jetson [12, 18, 19, 20, 22, 73, 13, 60, 68].
For platforms not requiring high computational performance on the vehicles themselves, microcontrollers [5, 32, 29, 64, 17, 70, 67], such as Arduino, are commonly employed. Although microcontrollers offer lower performance than Raspberry Pi, they enable real-time code execution, crucial for interfacing with low-level sensors and actuators. They also use less power and are cheaper.
Conversely, some platforms leverage regular PCs as their computational unit [10, 9, 262]. While PCs provide the highest performance, they have larger physical footprints and higher power consumption.

### v2

@@ IV-C5 Racing
IV-C6 Other tasks
In addition to the control tasks discussed earlier, we also explore tasks that are less commonly studied for small-scale cars, broadening the scope of research in this domain.
Drifting
Autonomous drifting, often associated with high-performance racing, represents a complex and sophisticated challenge in autonomous vehicle control. It requires precise coordination of speed, steering, throttle, and braking, alongside an in-depth understanding of vehicle dynamics and kinematics. Effective state estimation methods are equally critical to ensure stability and accuracy during drifting maneuvers. Traditional approaches heavily rely on dynamic models, such as the dynamic bicycle model used in [13], which employs EKF to fuse sensor data and integrates a Linear-Quadratic Regulator (LQR) controller with equilibrium drifting points for control. While this method is effective within its design constraints, its performance is tightly coupled to the accuracy of the dynamic model and the predefined operating conditions. Similarly, [218] expands upon this by introducing a six-state bicycle model, yet it struggles with generalizing to more complex and dynamic scenarios. This highlights a fundamental challenge: precise vehicle dynamics models are often impractical or unattainable in real-world environments. To address these limitations, [219] proposes an approach to tackle the drifting park problem with BARC cars by segmenting it into distinct phases: the normal driving regime and the sliding regime. During the normal driving phase, a nonlinear MPC operates with a predefined kinematic model. As the vehicle transitions into the sliding phase, reliance on this model diminishes. To navigate this shift, a feedforward-feedback controller takes charge, orchestrating safer maneuvers adeptly under sliding conditions. While effective, this segmentation introduces additional computational complexity and coordination challenges. Moreover, tire model singularities at low speeds exacerbate control difficulties, especially in the sliding regime. In response to these challenges, [220] presents a novel solution by integrating a fused kinematic-dynamic bicycle model with a nonlinear MPC framework tailored for a RACECAR platform. This unified approach harmonizes the planning and execution of dynamic vehicle maneuvers within a unified framework, optimizing the entire process cohesively. Despite these advancements, conventional dynamic models are constrained by their reliance on simplified assumptions, limiting their applicability. Future research should prioritize integrating data-driven methods, such as neural network-based controllers or DRL controllers, to improve adaptability and performance in uncertain environments. Additionally, the development of robust state estimation techniques that can operate effectively with noisy sensor inputs and less precise dynamic models is crucial for advancing autonomous drifting capabilities.
Parking
Besides driving behaviors, parking is also a crucial research area in AD for both small-scale and normal-scale cars. Research on the car parking problem generally stems from the broader motion planning problem and is typically defined as finding a collision-free path that connects the initial configuration to the final one. Traditional parking methods generally follow a three-phase approach: mapping the parking space, planning a collision-free path, and executing the maneuver. For small-scale cars, fuzzy logic algorithms dominate traditional methods. These systems rely on distance sensors, such as ultrasonic, infrared [224, 225, 226], or LiDAR [227, 228, 229, 230], to detect parking spaces and use predefined fuzzy rules to determine steering angles. While these methods are simple and effective in structured environments, they lack adaptability and struggle with complex, dynamic scenarios. Closed-loop controllers and rule-based strategies, while robust for specific cases, also suffer from limited flexibility [221, 222, 223].
ML-based methods offer adaptability and learning capabilities, making them better suited for dynamic parking environments. Neuro-fuzzy systems combine neural networks with fuzzy logic [250] to dynamically update rules and membership functions based on training data [231, 232], improving performance in diverse scenarios. Other ML approaches segment the parking process for better control, such as the use of a General Radial Basis Function (GRBF) classifier and Random Forest kernel to identify behavior transitions, as seen in [233]. Later, [234] proposed a two-stage learning framework to predict steering angles and gear status for parking using front and back-mounted monocular cameras. In the first stage, an encoder-decoder architecture estimates the initial steering angle trajectory. This trajectory, along with the heading angle and absolute position, is then fed into an LSTM network to estimate the optimal steering angle and gear status for parking. DRL approaches such as DDPG and other end-to-end learning frameworks have also shown promise in learning complex parking maneuvers, incorporating environmental feedback for precise control [235, 236]. While ML methods excel in adaptability, they often require large datasets, extensive training, and computational resources.
Despite the importance of autonomous parking for small-scale cars, research in this area has been lacking over the past few decades. However, with the development of current ML methods, there is a growing need to employ more advanced techniques in this research topic.
Cooperative Driving
Compared to the previously discussed scenarios, which typically involve a single controlled vehicle, cooperative driving entails a collaborative approach where multiple autonomous vehicles communicate, with V2V communications. In this setting, vehicles interact to achieve shared goals, navigate complex environments, and enhance traffic efficiency and safety. This involves exchanging data such as positions, speeds, and intended trajectories to enable collective decision-making, which is generally categorized into centralized and decentralized control strategies.
In centralized control, a central entity coordinates the actions of all vehicles by processing shared data to make globally optimized decisions. This approach excels in tasks like collaborative SLAM and multi-vehicle path planning, where precise coordination is crucial. For instance, [239] utilizes feature-based map registration for collaborative SLAM among three vehicles via V2V communication, enhancing mapping speed and accuracy. The centralized strategy excels in achieving global optimization and consistency but is limited by scalability challenges and potential single points of failure. The dependency on robust communication infrastructure also makes it vulnerable to latency and data synchronization issues. In contrast, decentralized control distributes decision-making among vehicles, which act based on local sensor data and peer communication. An example is the decentralized MPC method in [240], where MuSHR cars plan collision-free trajectories through negotiation, enabling precise and independent control. Safety is further addressed by [241], which proposes a priority-based distributed MPC (P-DMPC) algorithm to minimize collision risks. The decentralized control is advantageous for scalability and fault tolerance but can face challenges in achieving global optimization due to limited information sharing and potential inconsistencies among individual vehicle actions.
Specific applications of cooperative driving highlight its versatility. In roundabout navigation, vehicles communicate and negotiate entry, adjusting speeds and trajectories to yield to those already in the roundabout. RL-based coordination methods like those used in [58] and decentralized control frameworks from [56] demonstrate smooth merges and reduced stop-and-go traffic in scaled testbeds. Similarly, traffic jam mitigation emphasizes synchronized vehicle actions to maintain optimal flow, as seen in frameworks like [63], which adapts the IDM and MOBIL models for cooperative behavior, and studies like [242] and [243], which reduce travel times through decentralized optimal control. Emerging areas extend beyond autonomous cars. For example, [28] achieves adaptive cruise control for multi-vehicle platooning, while works such as [238] explore cooperation between UGVs and UAVs, pushing the boundaries of collaborative traffic management.
For future works, we propose to include hybrid control systems that combine centralized coordination with decentralized adaptability, robust V2V communication systems that minimize latency and security risks, and cross-platform collaboration between diverse vehicle types to enhance traffic management and operational flexibility. These advancements promise to unlock the full potential of cooperative driving in real-world applications.
IV-D End-to-end Driving
After discussing the modular system, we then delve into the end-to-end systems for small-scale cars. Here, we primarily focus on two main streams: lane keeping and racing.
For the lane keeping task, the end-to-end system has received more extensive research compared to modular systems. In this paradigm, the agent processes raw sensor input and outputs control commands directly, bypassing the modular perception, planning, and control layers. This framework primarily encompasses two methods: IL and DRL. Among IL methods, BC is foundational due to its simplicity, as it maps observed states directly to corresponding actions without modeling the underlying decision-making process. For example, in [48, 251], CNNs are trained with image data to predict directly the steering angle for control under various evaluation metrics within the DeepRacer platform. BC offers the advantages of straightforward implementation and computational efficiency, making it effective in scenarios where expert behavior is well-defined and consistent. However, its drawbacks include vulnerability to compounding errors, where deviations from expert behavior escalate in unseen states, and a heavy dependence on the quality and diversity of training data, which often hinders its ability to generalize to out-of-distribution scenarios. To address these issues, enhancements to BC have been proposed. For instance, [83] introduces uncertainty modeling during training to improve Sim2Real transfer on the Donkeycar platform. Similarly, [252] augments datasets using image style transfer [253] to enhance generalization. Comparative studies, such as [254], evaluate BC, GAIL, and DAgger in the Duckietown environment, highlighting the potential of Domain Adaptation techniques for better Sim2Real transfer. IDRL, in contrast, learns control policies directly from raw sensor data. Studies such as [10, 255, 256] employ raw images as input to DRL controllers, while preprocessing methods like dimensionality reduction are used in [257, 258, 90] to improve convergence speed. Addressing control smoothness, [259] introduces Conditioning for Action Policy Smoothness (CAPS), which reduces jerky control behaviors. Sim2Real transfer capabilities are improved by including delays and sampling rate as additional observations during training, as demonstrated by [37], enhancing the robustness of DRL policies on the DeepRacer platform. Beyond image inputs, raw LiDAR data serves as input in studies such as [46], where DRL algorithms navigate F1TENTH cars in structured environments. Although DRL eliminates the need for expert demonstrations and is highly adaptable to complex and dynamic environments, it often suffers from high computational cost and data inefficiency. Additionally, ensuring smooth and stable control remains challenging, particularly in real-world scenarios, and achieving reliable Sim2Real transfer is difficult due to discrepancies (Sim2Real gap) between training and deployment environments. To further enhance generalization and convergence, autoencoders are frequently integrated into end-to-end systems [260]. By compressing input data into lower-dimensional latent representations, autoencoders provide an efficient representation for downstream tasks such as control or policy learning. For example, [205, 206] preprocess image data with autoencoders before training Duckiebots using BC with expert trajectories. Variational Autoencoders (VAEs) [261] have also been explored in DRL, as demonstrated in [41, 208], where compressed latent features facilitate faster convergence and improved learning efficiency compared to raw image inputs. Despite their advantages in computational efficiency and feature extraction, autoencoders risk losing critical information during compression, particularly in dynamic or noisy environments, and introduce additional computational overhead during training and integration.
For racing with end-to-end systems, IL and DRL [262] have been widely explored. The authors in [82] introduce an end-to-end IL system, where a learner network needs to imitate an expert. The expert fuses GPS and IMU for state estimation and uses an MPC as controller, the learner uses a DNN as control policy to map raw, high-dimensional observations to continuous steering and throttle commands. Building on this, [263] presents a deep imitative RL framework for end-to-end racing with camera input, where IL is used to initialize the policy, and model-based RL is used for further refinement by interacting with an uncertainty-aware world model. This hybrid approach achieves improved adaptability and performance by leveraging the strengths of both IL and RL. To improve the stability of the racing system, [264] designs a residual control system, in which multiple Bayesian Neural Networks (BNNs) are trained with two camera inputs and GPS measurements to control an Autorally car in an end-to-end fashion. The integration of LiDAR technology has further expanded the capabilities of end-to-end systems. In [265], a model-based RL approach effectively utilizes raw LiDAR input to navigate racetracks with an F1TENTH car. Following this, [266] employs LiDAR data as input for a DQN to control the F1TENTH car, offering a comparative analysis of neural network architectures for LiDAR data processing. This study also evaluates two Sim2Real approaches, demonstrating the importance of transferability for real-world deployment. Despite the demonstrated potential of DRL for racing tasks, the well-known issue of low sample efficiency continues to hinder its broader application. For this issue, [267] propose an efficient residual policy learning method with the raw observation of LiDAR and IMU, in which first a controller based on the modified artificial potential field (MAPF) is used to generate policies, then DRL algorithms are used to generate a residual policy as a supplement to obtain the optimal policy with increased efficiency. Similarly, [268] presents a trajectory-aided learning (TAL) method that trains DRL with raw LiDAR input by incorporating the optimal trajectory into the learning formulation. [269] also presents a residual vehicle controller that learns to amend a traditional controller with a similar idea. To tackle the safety issues that usually occur in RL training, [270] uses a Viability Theory-based supervisor to recursively feasible vehicle safety during the training. Furthermore, to improve the robustness of RL, [271] first train a teacher model that overfits the training track, moving along a near-optimal path, then use this model to teach a student PPO model the correct actions along with randomization.
Looking ahead, hybrid approaches that combine IL and DRL for end-to-end systems show significant promise. Pretraining policies with IL to capture expert knowledge, followed by DRL fine-tuning for adaptability, can improve both sample efficiency and performance. Furthermore, advanced Sim2Real techniques, such as generative models and adversarial training, are critical for bridging the gap between simulated and real-world environments. Exploring multi-modal input fusion, and integrating diverse sensors like images, LiDAR, and radar, could enhance robustness and adaptability. Additionally, incorporating interpretable ML techniques would make the decision-making process of end-to-end systems more transparent, a crucial step for safety-critical applications. Finally, developing real-time adaptation algorithms that allow models to learn and adjust online without requiring retraining could ensure more reliable performance in dynamic, real-world settings. By addressing these challenges and pursuing these research directions, end-to-end systems can become more robust, efficient, and scalable, enabling broader adoption in real-world autonomous driving scenarios.
IV-E Proposed Framework
After reviewing the benchmarked tasks and techniques, it is evident that ML and DL methods form the backbone of AD research for small-scale cars. These methods play an important role across key modules such as perception, planning, and control. In perception, techniques like CNNs and autoencoders enhance object detection, lane detection, and semantic segmentation. For control, IL and RL enable models to make efficient and safe decisions in diverse and dynamic conditions. Connected driving leverages the Internet of Things (IoT) to create an interactive ecosystem, enabling real-time data exchange via V2V and V2I communication. This supports coordinated behaviors like platooning and optimized traffic flow management.
We would like to propose a baseline framework that integrates these methods into a cohesive structure. By extracting an effective framework from the literature, we aim to provide valuable insights for researchers. In a modular pipeline system, the process begins with the perception layer. Sensor fusion combines data from multiple sensors such as cameras and LiDAR using DNN, IMU, GPS, and encoders using KF alongside SLAM techniques. This integration ensures robust environmental perception and creates a comprehensive understanding of the environment. The fused feature or raw sensor input is then used for object detection, employing ML or shallow learning models to identify and classify objects in the vehicle’s vicinity. Compared with deep ML methods, shallow models have simple structures, usually consisting of one or a few layers of processing units, but are effective for resource-constrained environments, in our case, small-scale cars with lower-end computing units. Support vector machines or K-Nearest Neighbors can be used for perceptions. Subsequently, path planning and behavior planning are performed, either separately or within a single module. Classical methods like A*, Dijkstra’s, and RRT can be utilized for path planning, while behavior planning can leverage finite state machines, DRL, IL, MPC, and decision trees to handle tasks such as lane keeping, lane changing, overtaking, and obstacle avoidance. A safety and redundancy module ensures system reliability, incorporating PID or rule-based controllers for safe stopping during failures and redundant systems for robustness. For end-to-end systems, raw sensor data is processed directly using IL and RL or their hybrid approaches for driving tasks, with safety modules remaining integral. This cohesive framework aims to guide researchers in advancing small-scale autonomous driving systems.